Questions:

1. Compute the following ratios for PAYPAL HOLDINGS INC:

CURRENT RATIO

QUICK RATIO

CASH RATIO

TOTAL DEBT RATIO

DEBT EQUITY RATIO

TIMES INTEREST EARNED RATIO

CASH COVERAGE RATIO

INVENTORY TURNOVER

DAYS SALES IN INVENTORY

RECEIVABLES TURNOVER

DAYS SALES IN RECEIVABLES

TOTAL ASSET TURNOVER

CAPITAL INTENSITY

PROFIT MARGIN

RETURN ON ASSETS

RETURN ON EQUITY

PRICE EARNINGS RATIO

MARKET TO BOOK RATIO

2. Decompose the ROE using the extended Du-Pont Analysis.
Answer #1

1.

As of December 31,

2017

2016

(In millions, except par value)

ASSETS

Current assets:

Cash and cash equivalents

$

2,883

$

1,590

Short-term investments

2,812

3,385

Accounts receivable, net

283

214

Loans and interest receivable, net of allowances of $129 in 2017 and $339 in 2016

1,314

5,348

Loans and interest receivable, held for sale

6,398

Funds receivable and customer accounts

18,242

14,363

Prepaid expenses and other current assets

713

833

Total current assets

32,645

25,733

Long-term investments

1,961

1,539

Property and equipment, net

1,528

1,482

Goodwill

4,339

4,059

Intangible assets, net

168

211

Other assets

133

79

Total assets

$

40,774

$

33,103

LIABILITIES AND EQUITY

Current liabilities:

Accounts payable

$

257

$

192

Notes payable

1,000

Funds payable and amounts due to customers

19,742

15,163

Accrued expenses and other current liabilities

1,781

1,459

Income taxes payable

83

64

Total current liabilities

22,863

16,878

Deferred tax liability and other long-term liabilities

1,917

1,513

Total liabilities

24,780

18,391

Commitments and contingencies (Note 13)

Equity:

Common stock, $0.0001 par value; 4,000 shares authorized; 1,200 and 1,207 shares outstanding as of December 31, 2017 and 2016, respectively

Treasury stock at cost, 47 and 27 shares as of December 31, 2017 and 2016, respectively

(2,001

)

(995

)

Additional paid-in-capital

14,314

13,579

Retained earnings

3,823

2,069

Accumulated other comprehensive income (loss)

(142

)

59

Total equity

15,994

14,712

Total liabilities and equity

$

40,774

$

33,103

CONSOLIDATED STATEMENTS OF INCOME

Year Ended December 31,

2017

2016

2015

(In millions, except for per share amounts)

Net revenues

$

13,094

$

10,842

$

9,248

Operating expenses:

Transaction expense

4,419

3,346

2,610

Transaction and loan losses

1,011

1,088

809

Customer support and operations

1,364

1,267

1,110

Sales and marketing

1,128

969

937

Product development

953

834

792

General and administrative

1,155

1,028

873

Depreciation and amortization

805

724

608

Restructuring and other charges

132

48

Total operating expenses

10,967

9,256

7,787

Operating income

2,127

1,586

1,461

Other income (expense), net

73

45

27

Income before income taxes

2,200

1,631

1,488

Income tax expense

405

230

260

Net income

$

1,795

$

1,401

$

1,228

Net income per share:

Basic

$

1.49

$

1.16

$

1.00

Diluted

$

1.47

$

1.15

$

1.00

Weighted average shares:

Basic

1,203

1,210

1,222

Diluted

1,221

1,218

1,229

Shares Repurchased

Average Price
Paid per Share
(1)

Value of Shares Repurchased

Remaining Amount Authorized for Repurchases

(In millions, except per share amounts)

Period ended October 31, 2017

$

5,299

Period ended November 30, 2017

$

5,299

Period ended December 31, 2017

4.0

$

74.30

$

300

$

4,999

4.0

$

300

Current Ratio = Current Assets / Current Liabilities = 32645 / 22863 = 1.428

Quick Ratio = (Current Assets – Inventory) / Current Liabilities = (32645- 0) / 22863 = 1.428

Cash Ratio = Cash / Current Liabilities = 2883 / 22863 = 0.126

Total Debt Ratio = Total Debt / Total Assets = 24780 / 40774 = 0.607

Debt Equity Ratio = Total Debt / Total Equity = 24780 / 15994 = 1.549

Times Interest Earned Ratio = Earnings Before Interest & Tax / Interest Payments   = 2127 / 0 = Not Applicable

Cash Coverage Ratio = (Earnings Before Interest & Tax + Depreciation) / Interest Payments = (2127+805) / 0 = Not Applicable

Inventory Turnover = Cost of Goods Sold / Average Inventory   = 0    [As, COGS is 0]

Days Sales In Inventory = 365 / Inventory Turnover = Not Applicable

Receivables Turnover = Sales / Average Receivables = 13094 / {(283+214)/2} = 52.69

[Average Receivables = (Ending Receivables + Opening Receivables) / 2]

Days Sales In Receivables = 365 / Receivables Turnover = 365 / 52.69 = 6.927

Total Asset Turnover = Sales / Total Assets = 13094 / 40774 = 0.321

Capital Intensity = Total Assets / Sales = 40774 / 13094 = 3.114

Profit Margin = Net income / Sales = 1795 / 13094 = 0.137

Return On Assets = Net income / Total Assets = 1795 / 40774 = 0.044

Return On Equity = Net income/ Total Equity = 1795 / 15994 = 0.112

Price Earnings Ratio = Price per Share / Earning per share = 74.30 / 1.49 = 49.86

Market To Book Ratio = Market value per share / Book value per share = 74.30/ (15994/1203)   = 5.589

[Book value per share = Total Equity / No. of shares outstanding]

2. Du-Pont Analysis:

Return On Equity (ROE) = Net income/ Total Equity = (Net income / Sales) x (Sales / Total Assets) x (Total Assets / Total Equity)

= (1795 / 13094) x (13094 / 40774) x (40774 /15994) = 0.112